Suppose you know that ∠S and ∠Y are complementary, and that m∠S = 3(m∠Y) − 150°. Find m∠Y
Answer by math_tutor2020(3816)


Answer: 60 degrees

---------------------------------------------------

Reasoning:

Complementary angles add to 90 degrees by definition.
Complementary angles form a corner.
Both "complementary" and "corner" start with "Co", and this is the trick you can use to remember the definition.

In contrast, supplementary angles add to 180 degrees
The angles form a straight line.
"supplementary" and "straight" both start with S.

Anyways, the angles S and Y add to 90 degrees, which means,
(angle S) + (angle Y) = 90
(3Y-150) + (Y) = 90
4Y - 150 = 90
4Y = 90+150
4Y = 240
Y = 240/4
Y = 60
Angle Y is 60 degrees

From there, we can find angle S if you wanted
S = 3Y - 150
S = 3*60 - 150
S = 180 - 150
S = 30

As a way to check,
(angle S) + (angle Y) = 60 + 30 = 90
The answer is confirmed.